If a 25- sided fair die that is numbered from 1 to 25 is rolled once, find the probability of getting

(a) a number less than 5 or at least 20. Show your work in details.

(b) a number less than 5 and at least 20. Show your work in details

(c) an even number. Show your work in details

(d) an odd number. Show your work in details
Answer #1

(a)

Out of 25 numbers 4 are less than 5 so

P(less than 5) = 4/25

Out of 25 numbers, 6 are 20 or greater than 20 so

P(at least 20) = 6/20

Since above events are mutually exclusive so

P(less than 5 or at least 20) = 4/25 + 6/25 = 10/25 = 0.40

(b)

Since events are mutually exclusive so

P(less than 5 and at least 20) = 0

(c)

Out of 25 numbers, 12 are even so

P(even) = 12 / 25 = 0.48

(d)

Out of 25 numbers, 13 are odd so

P(odd) = 13 / 25 = 0.52